# Retinex-RAWMamba: Bridging Demosaicing and Denoising for Low-Light RAW Image Enhancement

### v1

@@ Abstract
Low-light image enhancement, particularly in cross-domain tasks such as mapping from the raw domain to the sRGB domain, remains a significant challenge. Many deep learning-based methods have been developed to address this issue and have shown promising results in recent years. However, single-stage methods, which attempt to unify the complex mapping across both domains, leading to limited denoising performance. In contrast, two-stage approaches typically decompose a raw image with color filter arrays (CFA) into a four-channel RGGB format before feeding it into a neural network. However, this strategy overlooks the critical role of demosaicing within the Image Signal Processing (ISP) pipeline, leading to color distortions under varying lighting conditions, especially in low-light scenarios. To address these issues, we design a novel Mamba scanning mechanism, called RAWMamba, to effectively handle raw images with different CFAs. Furthermore, we present a Retinex Decomposition Module (RDM) grounded in Retinex prior, which decouples illumination from reflectance to facilitate more effective denoising and automatic non-linear exposure correction. By bridging demosaicing and denoising, better raw image enhancement is achieved. Experimental evaluations conducted on public datasets SID and MCR demonstrate that our proposed RAWMamba achieves state-of-the-art performance on cross-domain mapping.
Codehttps://github.com/Cynicarlos/RetinexRawMamba
Introduction
Existing deep learning methods, particularly those focused on low-light enhancement tasks, primarily operate in the sRGB domain. However, RAW images typically possess a higher bit depth than their RGB counterparts, meaning they retain a greater amount of original detail. Consequently, processing from RAW to RGB is often more effective. However, RAW and RGB are distinct domains with image processing algorithms tailored to their specific characteristics. For instance, in the RAW domain, algorithms prioritize denoising, whereas in the RGB domain, they focus on color correction. This difference often renders single-stage end-to-end methods (Maharjan et al. 2019; Gu et al. 2019; Lamba and Mitra 2021) ineffective.
@@ Introduction
Demosaicing algorithms play a crucial role in converting RAW image to sRGB, with most traditional methods relying on proximity interpolation. Although some researchers have explored CNN-based approaches (Xu et al. 2020; Dong et al. 2022) to map noisy RAW images to clean sRGB outputs, the limited receptive field inherent in convolutional networks often hampers their effectiveness in demosaicing tasks. To address this, Vision Transformers (ViTs) have been employed to expand the receptive field, but the attention mechanisms in ViTs are computationally intensive. The introduction of Mamba provides a more efficient balance between these trade-offs. However, existing Mamba scanning mechanisms do not adequately address the diverse characteristics of RAW images with different Color Filter Arrays (CFAs), highlighting the need for Mamba scanning methods specifically tailored to various CFAs.
@@
Additionally, previous methods (Chen et al. 2018; Jin et al. 2023a) for processing short-exposure RAW images often rely on a simple linear multiplication of a prior for exposure correction. Specifically, short-exposure RAW images, which contain significant noise, are multiplied by the exposure time ratio of the corresponding long-exposure image. This approach assumes uniform exposure across the image, which is often unrealistic and can result in sub-optimal denoising and inaccurate brightness. By leveraging the success of the Retinex theory in low-light enhancement tasks for RGB images (Land and McCann 1971; Ma et al. 2022; Sun et al. 2024a), we introduce a Retinex-based dual-domain auxiliary exposure correction method, namely Retinex Decomposition Module (RDM), which decouples illumination and reflection and realize automatic nonlinear exposure correction to achieve more efficient denoising effect and more accurate brightness correction. Furthermore, given the significant differences in noise distribution between different RAW domain and sRGB domain, we build upon the idea of decoupling the task into two sub-tasks: denoising on the raw domain (Yi et al. 2021; Jin et al. 2023b; Wang et al. 2020) and cross-domain mapping (Schwartz, Giryes, and Bronstein 2018; Zamir et al. 2020; Sun et al. 2024b; Ignatov, Van Gool, and Timofte 2020).
In general, we propose a Retinex-based decoupling network (Retinex-RAWMamba) for RAW domain denoising and low-light enhancement shown in Fig. 2. RAWMamba decouples the tasks of denoising and demosaicing into two distinct sub-tasks, effectively mapping noisy RAW images to clean sRGB images. Specifically, for demosaicing sub-task, we introduce RAWMamba to fully consider all pixels in the immediate neighborhood of a certain pixe by utilizing eight direction mechanism. For the denoising sub-task, we propose the Retinex Decomposition Module, which enhances both denoising performance and brightness correction.
@@
Figure 2: The overall architecture of our proposed Retinex-RAWMamba and (a) Retinex Decomposition Module, (b) Domain Adaptive Fusion and (c) Simple Denoising Block
Related Work
Low Light Enhancement on Raw Domain
@@
In the Raw domain low-light enhancement task, researchers have proposed some innovative approaches. Since the task can be split into two sub-tasks, RAW domain denoising and color correction, some of the work only focuses on one of the sub-tasks. For example, on the raw domain denoising task, there are noise modeling with deep learning methods (Bao et al. 2020; Cao et al. 2023; Feng et al. 2024a, b; Wei et al. 2022; Zou and Fu 2022), which ultimately compute evaluation metrics on the raw domain. After the release of the SID public dataset by Chen at al.(Chen et al. 2018) 2018, researchers have proposed many works that address both tasks simultaneously. These works can be further categorized into single-stage approaches and multi-stage approaches. Single-stage methods (Chen et al. 2018; Zamir et al. 2019) aims to map noisy raw to clean sRGB by training a single model. For instance, SID (Chen et al. 2018) only used a simple UNet to accomplish this task. DID (Maharjan et al. 2019) proposed a deep neural network based on residual learning for end-to-end extreme low-light image denoising. SGN (Gu et al. 2019) introduced a self-guided network, which adopted a top-down self-guidance architecture to better exploit image multi-scale information. Since the ISP undergoes many nonlinear transformations, it is still difficult to learn for a single neural network, and it can only be realized by piling up a large number of parameters, which leads to inefficiencies, and thus multi-stage methods came into being. Multi-stage methods (Huang et al. 2022; Xu et al. 2020; Zhu et al. 2020) achieve better results by decoupling the tasks, this idea effectively reduces the ambiguity between different domains. For instance, Huang et al. (Huang et al. 2022) proposed intermediate supervision on the raw domain, while Dong et al. (Dong et al. 2022) did that on the monochrome domain. DNF (Jin et al. 2023a) introduced a decoupled two-stage net with weight weight-shared encoder to reduce the number of parameters while achieving good results.
Mamba in Vision Task
@@
State Space Models (SSM) are recently introduced to deep learning since they can effectively model long range dependencies. For instance, (Gu, Goel, and Ré 2022) proposes a Structured State-Space Sequence (S4) model and recently, (Gu and Dao 2023) proposes Mamba, which outperforms Transformers at various sizes on large-scale real data and enjoys linear scaling in sequence length. In addition to Mamba’s great work on NLP tasks, researchers have also made many attempts and achieved good results on visual tasks, such as classification (Chen et al. 2024; Yue and Li 2024), segmentation (Ma, Li, and Wang 2024; Wang et al. 2024; Xing et al. 2024; Ruan and Xiang 2024; Liu et al. 2024a), generation (Shen et al. 2024; Hu et al. 2024), and image restoration (Guo et al. 2024; Shi et al. 2024; Zhen, Hu, and Feng 2024; Zheng and Wu 2024; Pei, Huang, and Xu 2024; Bai et al. 2024). EfficientVMamba (Pei, Huang, and Xu 2024) presents the Efficient 2D Scanning (ES2D) method, utilizing atrous sampling of patches on the feature map to speed up training. VMamba (Liu et al. 2024b) incorporates a Cross-Scan Module (CSM), which converts the input image into sequences of patches along the horizontal and vertical axes, and it enables the scanning of sequences in four distinct directions. That is, each pixel integrates information from the four surrounding pixels. FreqMamba (Zhen, Hu, and Feng 2024) introduces complementary triple interaction structures including spatial Mamba, frequency band Mamba, and Fourier global modeling, which utilizes the complementarity between Mamba and frequency analysis for image deraining.
Method
Figure 3: Details of (a) RAWMamba, (b) Residual RAWMamba and (c) SS2D
Preliminaries
State Space Model (SSM)
@@
The overall pipeline is shown in Fig. 2. First, we preprocess the low-exposure noisy single-channel raw image by multiplying it with the exposure time ratio of the long-exposure ground truth (GT). Then, based on the Color Filter Array (CFA) pattern, we pack it into a multi-channels input. Specifically, for Bayer format, we pack the input $\textbf{X}\in\mathbb{R}^{H\times W\times 1}$ into four channels input $\textbf{X}_{packed}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times 4}$; for XTrans format, we pack the input into nine channels input $\textbf{X}_{packed}\in\mathbb{R}^{\frac{H}{3}\times\frac{W}{3}\times 9}$. Both stages of Retinex-RAWMamba are built upon the UNet-based (Ronneberger, Fischer, and Brox 2015) encoder-decoder architecture. The first stage of the overall framework is dedicated to raw domain denoising. Initially, The Retinex Decomposition Module (RDM) processes the input to generate both the primary input and an auxiliary illumination feature map, denoted as lf. This primary input is then fused with $\textbf{{lf}}_{i}$ at each encoding layer and denoised using the Simple Denoising Block (SDB), which consists of three consecutive convolutions followed by a residual connection. At each decoding layer, a temporary feature map $\textbf{{ft}}_{i}$ is generated, ultimately resulting in the generation of the denoised raw image. The loss function for the first stage is computed against the ground truth RAW image, which also serves as the supervision signal for this stage. In the second stage, the focus shifts to demosaicing and color correction. In this phase, the input is fused with both $\textbf{{lf}}_{i}$ and $\textbf{{ft}}_{i}$, and processed by the RAWMamba at each layer. After decoding, the final RGB image is produced. The loss function is then calculated against the ground truth RGB image, providing the supervision signal for this stage and guiding the optimization of whole model.
RAWMamba
@@
The details of RAWMamba and Residual RAWMamba are shown in Fig. 3. RAWMamba leverages the naive visual mamba in MambaIR (Guo et al. 2024), with an innovative scanning mechanism. In the ISP process from Raw to RGB, proximity interpolation is commonly employed for demosaicing and often involves considering all eight closely connected locations around a given position, and Fig.1 (a) gives an example with Bayer pattern raw image, (b) shows the scanning in RAWMamba (black dashed line) and naive Mamba (purple dashed line). The naive scanning method fails to consider the continuity of scanning, resulting in a lack of continuity between the end of each row/column and its bottom/right side. This leads to gaps in image semantics, which hinders image reconstruction. To address this issue, we propose using a Z-scan. That is, when the scan reaches the end of each row/column, the reverse scan starts from the next row/column immediately adjacent to the last pixel. However, there are still limitations with this scanning method as it does not take into account all eight surrounding pixels when certain pixels are close to each other at the top, bottom, left, and right positions. Taking into consideration the characteristics of this task, we introduce Eight direction Mamba.
@@
The detail of the our proposed scan mechanism is shown in 3 (c). Specifically, for a feature map $\textbf{{F}}\in{\mathbb{R}}^{C\times H\times W}$, we first flip its even rows (fer) and columns (fec) to get $\textbf{{F}}_{fer}\in{\mathbb{R}}^{C\times H\times W}$ and $\textbf{{F}}_{fec}\in{\mathbb{R}}^{C\times H\times W}$, respectively. Then we flatten $\textbf{{F}}_{fer}$ and $\textbf{{F}}_{fec}$ to get the first tow directions’ scanning $\textbf{{F}}_{1}\in{\mathbb{R}}^{C\times HW}$ and $\textbf{{F}}_{2}\in{\mathbb{R}}^{C\times HW}$. Then we can get the feature of the oblique scan $\textbf{{F}}_{3}$ as following:
@@
$RAWMamba(\textbf{{F}})=Reshape(\sum_{i=1}^{8}\overline{\textbf{{F}}}_{i},(C,H,W))$ (6)
For the proposed Residual RAWMamba, It can be simply formulated as:
T $=\alpha\textbf{{X}}+RAWMamba(LN(\textbf{{X}})))$ (7)
$Out$ $=\beta\textbf{{T}}+CA(GELU(Conv(LN(\textbf{{T}}))))$
@@
where, LN is layer normalization, $\alpha$ and $\beta$ are parameters that can be learned, CA is channel attention.
Retinex Decomposition
@@
Low-light enhancement methods based on retinex theory have been successful in RGB domain (Wei et al. 2018; Wu et al. 2022; Cai et al. 2023), so we propose dual-domain Retinex Decomposition Module. RDM can decompose image $\textbf{{X}}\in{\mathbb{R}}^{H\times W\times C}$ into the reflection component $\textbf{{R}}\in{\mathbb{R}}^{H\times W\times C}$ and the illumination component $\textbf{{I}}\in{\mathbb{R}}^{H\times W\times C}$. The details of RDM are shown in Fig. 2 (a). The module first takes the average value of the input image $\textbf{{X}}\in{\mathbb{R}}^{H\times W\times C}$ in the channel dimension to obtain $\textbf{{M}}\in{\mathbb{R}}^{H\times W\times 1}$, concatenates them in the channel dimension, and then passes several convolutions to obtain the first output light feature $\textbf{{I}}\in{\mathbb{R}}^{H\times W\times C}$, and then passes a $1\times 1$ convolution to obtain the light map, and finally light map will be multiplied by the original input X to obtain the second output reflection map R after a $1\times 1$ convolution. Specifically,
I $=Convs_{1,5,3}\{cat[\textbf{{X}},sap(\textbf{{X}})]\}$ (8)
R $=\textbf{{X}}\odot Conv_{1}(\textbf{{I}})$
where sap donates Spatial Average Pooling, cat refers to the concatenation of tow feature maps on the channel dimension, $Convs_{1,5,3}$ donates a series of convolution with kernel size 1, 5 and 3, $\odot$ is the Hadamard product. And after obtaining the light feature I, we will simply downsample it to get four feature maps at each layer, which are donated as $\textbf{{lf}}_{i},i=1,2,3,4\}$, $\textbf{{lf}}_{i}$ is the $i_{th}$ layer light feature that will be fused later for auxiliary automatic exposure correction at layer $i$.
Table 1: Quantitative results of RAW-based LLIE methods on the Sony and Fuji subsets of SID. The top-performing result is highlighted in bold, while the second-best is shown in underline. Metrics marked with $\uparrow$ indicate that a higher value is better, and those marked with $\downarrow$ indicate that a lower value is better. N/A indicates the result is not available.
Category Method Venue #Params.(M) Sony Fuji
Single-Stage SID CVPR2018 7.7 28.96 0.787 0.356 26.66 0.709 0.432
DID ICME2019 2.5 29.16 0.785 0.368 N/A N/A N/A
SGN ICCV2019 19.2 29.28 0.790 0.370 27.41 0.720 0.430
LLPackNet BMVC2020 1.2 27.83 0.755 0.541 N/A N/A N/A
RRT CVPR2021 0.8 28.66 0.790 0.397 26.94 0.712 0.446
Multi-Stage EEMEFN AAAI2020 40.7 29.60 0.795 0.350 27.38 0.723 0.414
LDC CVPR2020 8.6 29.56 0.799 0.359 27.18 0.703 0.446
MCR CVPR2022 15.0 29.65 0.797 0.348 N/A N/A N/A
RRENet TIP2022 15.5 29.17 0.792 0.360 27.29 0.720 0.421
DNF CVPR2023 2.8 30.62 0.797 0.343 28.71 0.726 0.391
Ours - 6.2 30.76 0.810 0.328 29.02 0.743 0.382
Domain Adaptive Fusion

### v2

@@ Abstract
Low-light image enhancement, particularly in cross-domain tasks such as mapping from the raw domain to the sRGB domain, remains a significant challenge. Many deep learning-based methods have been developed to address this issue and have shown promising results in recent years. However, single-stage methods, which attempt to unify the complex mapping across both domains, leading to limited denoising performance. In contrast, two-stage approaches typically decompose a raw image with color filter arrays (CFA) into a four-channel RGGB format before feeding it into a neural network. However, this strategy overlooks the critical role of demosaicing within the Image Signal Processing (ISP) pipeline, leading to color distortions under varying lighting conditions, especially in low-light scenarios. To address these issues, we design a novel Mamba scanning mechanism, called RAWMamba, to effectively handle raw images with different CFAs. Furthermore, we present a Retinex Decomposition Module (RDM) grounded in Retinex prior, which decouples illumination from reflectance to facilitate more effective denoising and automatic non-linear exposure correction. By bridging demosaicing and denoising, better raw image enhancement is achieved. Experimental evaluations conducted on public datasets SID and MCR demonstrate that our proposed RAWMamba achieves state-of-the-art performance on cross-domain mapping. The code is available at https://github.com/Cynicarlos/RetinexRawMamba.
1 Introduction
Existing deep learning methods, particularly those focused on low-light enhancement tasks, primarily operate in the sRGB domain. However, RAW images typically possess a higher bit depth than their RGB counterparts, meaning they retain a greater amount of original detail. Consequently, processing from RAW to RGB is often more effective. However, RAW and RGB are distinct domains with image processing algorithms tailored to their specific characteristics. For instance, in the RAW domain, algorithms prioritize denoising, whereas in the RGB domain, they focus on color correction. This difference often renders single-stage end-to-end methods [26, 12, 19] ineffective.
@@ 1 Introduction
Demosaicing algorithms play a crucial role in converting RAW image to sRGB, with most traditional methods relying on proximity interpolation. Although some researchers have explored CNN-based approaches [42, 7] to map noisy RAW images to clean sRGB outputs, the limited receptive field inherent in convolutional networks often hampers their effectiveness in demosaicing tasks. To address this, Vision Transformers (ViTs) have been employed to expand the receptive field, but the attention mechanisms in ViTs are computationally intensive. The introduction of Mamba provides a more efficient balance between these trade-offs. However, existing Mamba scanning mechanisms do not adequately address the diverse characteristics of RAW images with different Color Filter Arrays (CFAs), highlighting the need for Mamba scanning methods specifically tailored to various CFAs.
@@
Additionally, previous methods [5, 17] for processing short-exposure RAW images often rely on a simple linear multiplication of a prior for exposure correction. Specifically, short-exposure RAW images, which contain significant noise, are multiplied by the exposure time ratio of the corresponding long-exposure image. This approach assumes uniform exposure across the image, which is often unrealistic and can result in sub-optimal denoising and inaccurate brightness. By leveraging the success of the Retinex theory in low-light enhancement tasks for RGB images [21, 25, 33], we introduce a Retinex-based dual-domain auxiliary exposure correction method, namely Retinex Decomposition Module (RDM), which decouples illumination and reflection and realize automatic nonlinear exposure correction to achieve more efficient denoising effect and more accurate brightness correction. Furthermore, given the significant differences in noise distribution between different RAW domain and sRGB domain, we build upon the idea of decoupling the task into two sub-tasks: denoising on the raw domain [43, 18, 35] and cross-domain mapping [30, 46, 34, 16].
In general, we propose a Retinex-based decoupling network (Retinex-RAWMamba) for RAW domain denoising and low-light enhancement shown in Fig. 2. Our method decouples the tasks of denoising and demosaicing into two distinct sub-tasks, effectively mapping noisy RAW images to clean sRGB images. Specifically, for demosaicing sub-task, we introduce RAWMamba to fully consider all pixels in the immediate neighborhood of a certain pixel by utilizing eight direction mechanism. For the denoising sub-task, we propose the Retinex Decomposition Module, which enhances both denoising performance and brightness correction. Additionally, we introduce a dual-domain encoding stage enhance branch designed to leverage the meticulously preserved detail features from the raw domain, thereby compensating for the information loss that occurs during the denoising phase.
@@
Figure 2: The overall architecture of our proposed Retinex-RAWMamba and (a) Retinex Decomposition Module, (b) Simple Denoising Block and (c) Domain Adaptive Fusion
2 Related Work
2.1 Low Light Enhancement on Raw Domain
@@
In the Raw domain low-light enhancement task, researchers have proposed some innovative approaches. Since the task can be split into two sub-tasks, RAW domain denoising and color correction, some of the work only focuses on one of the sub-tasks. For example, on the raw domain denoising task, there are noise modeling with deep learning methods [2, 4, 8, 9, 39, 51], which ultimately compute evaluation metrics on the raw domain. After the release of the SID public dataset by Chen at al.[5] 2018, researchers have proposed many works that address both tasks simultaneously. These works can be further categorized into single-stage approaches and multi-stage approaches. Single-stage methods [5, 45] aims to map noisy raw to clean sRGB by training a single model. For instance, SID [5] only used a simple UNet to accomplish this task. DID [26] proposed a deep neural network based on residual learning for end-to-end extreme low-light image denoising. SGN [12] introduced a self-guided network, which adopted a top-down self-guidance architecture to better exploit image multi-scale information. Since the ISP undergoes many nonlinear transformations, it is still difficult to learn for a single neural network, and it can only be realized by piling up a large number of parameters, which leads to inefficiencies, and thus multi-stage methods came into being. Multi-stage methods [15, 42, 50] achieve better results by decoupling the tasks, this idea effectively reduces the ambiguity between different domains. For instance, Huang et al. [15] proposed intermediate supervision on the raw domain, while Dong et al. [7] did that on the monochrome domain. DNF [17] introduced a decoupled two-stage net with weight weight-shared encoder to reduce the number of parameters while achieving good results.
2.2 Mamba in Vision Task
@@
State Space Models (SSM) are recently introduced to deep learning since they can effectively model long range dependencies. For instance, [11] proposes a Structured State-Space Sequence (S4) model and recently, [10] proposes Mamba, which outperforms Transformers at various sizes on large-scale real data and enjoys linear scaling in sequence length. In addition to Mamba’s great work on NLP tasks, researchers have also made many attempts and achieved good results on visual tasks, such as classification [6, 44], segmentation [24, 37, 41, 29, 22], generation [31, 14], and image restoration [13, 32, 48, 49, 27, 1]. EfficientVMamba [27] presents the Efficient 2D Scanning (ES2D) method, utilizing atrous sampling of patches on the feature map to speed up training. VMamba [23] incorporates a Cross-Scan Module (CSM), which converts the input image into sequences of patches along the horizontal and vertical axes, and it enables the scanning of sequences in four distinct directions. That is, each pixel integrates information from the four surrounding pixels. FreqMamba [48] introduces complementary triple interaction structures including spatial Mamba, frequency band Mamba, and Fourier global modeling, which utilizes the complementarity between Mamba and frequency analysis for image deraining.
3 Method
3.1 Preliminaries
3.1.1 State Space Model (SSM)
@@ 3.2 Overall Pipeline
The overall pipeline is shown in Fig. 2. First, we preprocess the low-exposure noisy single-channel raw image by multiplying it with the exposure time ratio of the long-exposure ground truth (GT). Then, based on the Color Filter Array (CFA) pattern, we pack it into a multi-channels input. Specifically, for Bayer format, we pack the input $\textbf{X}\in\mathbb{R}^{H\times W\times 1}$ into four channels input $\textbf{X}_{packed}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times 4}$; for XTrans format, we pack the input into nine channels input $\textbf{X}_{packed}\in\mathbb{R}^{\frac{H}{3}\times\frac{W}{3}\times 9}$. Both stages of Retinex-RAWMamba are built upon the UNet-based [28] encoder-decoder architecture. The first stage of the overall framework is dedicated to raw domain denoising. Initially, The Retinex Decomposition Module (RDM) processes the input to generate tow feature maps L and R, L will be multiplied by the original input to obtain $\textbf{X}_{in}$ and R will be used later. $\textbf{X}_{in}$ will pass the denoising stage and generate the first output $\textbf{O}_{1}\in\mathbb{R}^{H\times W\times C_{in}}$. Then $\textbf{X}_{in}$ will pass the demosaicing stage to generate the second output $\textbf{O}_{2}\in\mathbb{R}^{H\times W\times 3}$. The overall loss function is then calculated against both ground truth RAW and ground truth RGB images, providing the supervision signal for both domains and guiding the optimization of whole model.
Figure 3: Details of (a) RAWMamba, (b) RAWSSM and (c) SS2D
3.3 RAWMamba
@@
The details of RAWMamba and is shown in Fig. 3 (a). The RAWSSM leverages the naive visual mamba in MambaIR [13], with an innovative scanning mechanism. In the ISP process from Raw to RGB, proximity interpolation is commonly employed for demosaicing and often involves considering all eight closely connected locations around a given position, and Fig. 1 (a) gives an example with Bayer pattern raw image, (b) shows the scanning in RAWMamba (black dashed line) and naive Mamba (purple dashed line). The naive scanning method fails to consider the continuity of scanning, resulting in a lack of continuity between the end of each row/column and its bottom/right side. This leads to gaps in image semantics, which hinders image reconstruction. To address this issue, we propose using a Z-scan. That is when the scan reaches the end of each row/column, the reverse scan starts from the next row/column immediately adjacent to the last pixel. However, there are still limitations with this scanning method as it does not take into account all eight surrounding pixels when certain pixels are close to each other at the top, bottom, left, and right positions. Taking into consideration the characteristics of this task, we introduce Eight direction Mamba.
@@
The detail of the our proposed scan mechanism is shown in Fig. 3 (c). Specifically, for a feature map $\textbf{{F}}\in{\mathbb{R}}^{C\times H\times W}$, we first flip its even rows (fer) and columns (fec) to get $\textbf{{F}}_{fer}\in{\mathbb{R}}^{C\times H\times W}$ and $\textbf{{F}}_{fec}\in{\mathbb{R}}^{C\times H\times W}$, respectively. Then we flatten $\textbf{{F}}_{fer}$ and $\textbf{{F}}_{fec}$ to get the first two directions’ scanning $\textbf{{F}}_{1}\in{\mathbb{R}}^{C\times HW}$ and $\textbf{{F}}_{2}\in{\mathbb{R}}^{C\times HW}$. Then we can get the feature of the oblique scan $\textbf{{F}}_{3}$ as following:
@@
$SS2D(\textbf{{F}})=Reshape(\sum_{i=1}^{8}\overline{\textbf{{F}}}_{i},(C,H,W))$ (6)
For the RAWSSM, given an input X, it can be formulated as follows:
$x,z$ $=chunk(Linear(\textbf{{X}}))$ (7)
$x$ $=LN(SS2D(SiLU(Conv_{3}(x))))$
$out$ $=x*SiLU(z)$
where out is the output of RAWSSM, LN is layer normalization, $Conv_{3}$ is the convolution operation with a kernel size of $3\times 3$, SiLU is the activate function.
And for the proposed RAWMamba, given an input X, it can be simply formulated as:
$t$ $=\alpha\textbf{{X}}+RAWSSM(LN(\textbf{{X}})))$ (8)
$out$ $=\beta t+CA(GELU(Conv(LN(t))))$
@@
where, out is the output of RAWMamba, $\alpha$ and $\beta$ are parameters that can be learned, CA is channel attention.
3.4 Retinex Decomposition
@@
Low-light enhancement methods based on retinex theory have been successful in RGB domain [38, 40, 3], so we propose dual-domain Retinex Decomposition Module. RDM can decompose image $\textbf{{X}}\in{\mathbb{R}}^{H\times W\times C_{in}}$ into the reflection component $\textbf{{R}}\in{\mathbb{R}}^{H\times W\times C}$ and the illumination component $\textbf{{L}}\in{\mathbb{R}}^{H\times W\times C_{in}}$. The details of RDM are shown in Fig. 2 (a). The module first takes the average value of the input image $\textbf{{X}}\in{\mathbb{R}}^{H\times W\times C_{in}}$ in the channel dimension to obtain $\textbf{{M}}\in{\mathbb{R}}^{H\times W\times 1}$, concatenates it in the channel dimension, and then passes several convolutions and a GELU activate function to obtain the first output $\textbf{{R}}\in{\mathbb{R}}^{H\times W\times C}$, and then passes a $1\times 1$ convolution to obtain the light map L , which will be multiplied by the original input to pre-adjust the light. Specifically,
L $=Convs_{1,5,3}\{cat[\textbf{{X}},mean(\textbf{{X}},dim=-1)]\}$ (9)
R $=Conv_{1}(\textbf{{L}})$
where cat refers to the concatenation of two feature maps on the channel dimension, $Convs_{1,5,3}$ donates a series of convolution with kernel size 1, 5 and 3.
3.5 Dual-domain Encoding Stage Enhance Branch
Considering that the feature R obtained from the RDM contains most of the details that could be lost after the first stage, we make full use of these features in both domains and in order to reduce the amount of calculation, we propose dual-domain encoding stage enhance branch, which does not be used in the decoding stages. Specifically, after obtaining R, we will simply downsample it to get four feature maps at each layer, which are donated as $\textbf{{R}}_{i},i=1,2,3,4\}$, $\textbf{{R}}_{i}$ is the $i_{th}$ layer light feature that will be fused later with DAF for auxiliary automatic exposure correction at layer $i$. At the $i_{th}$ layer of the encoding stage during the denoising phase, the denoising feature $\textbf{{dn}}_{i}$ will firstly fused with $\textbf{{R}}_{i}$ before passing the SDB. Similarly, at the $i_{th}$ layer of the encoding stage during the demosaicing phase. the the demosaicing feature $\textbf{{dm}}_{i}$ will firstly fused with $\textbf{{R}}_{i}$ before passing the RAWMamba. This method of performing fusion only in the encoding stage fully utilizes the features that are not lost to improve the ability to restore details while reducing the amount of calculation.
3.6 Domain Adaptive Fusion